OpenStudy (pgreene008):

Find the angle between the given vectors to the nearest tenth of a degree. u = <8, 4>, v = <9, -9> 25.8° 35.8° 71.6° 81.6°

OpenStudy (pgreene008):

@wolf1728 can i get your help again?

OpenStudy (kainui):

This might help: \[\vec u \cdot \vec v = |u||v| \cos \theta \]

OpenStudy (pgreene008):

i still dont get it

OpenStudy (pgreene008):

25.8° 35.8° 71.6° 81.6° im supposed to get one of those

OpenStudy (pgreene008):

@TheSmartOne

OpenStudy (pgreene008):

@wio

OpenStudy (anonymous):

Do you know the dot product?

OpenStudy (pgreene008):

not really, no

OpenStudy (anonymous):

do you know how graph points?

OpenStudy (pgreene008):

no

OpenStudy (pgreene008):

@wio

OpenStudy (anonymous):

You can't do this problem if you're not even able do graph vectors and guess the angle between them.

OpenStudy (pgreene008):

i tried it and i got 71.6 degrees

OpenStudy (anonymous):

Okay, good

OpenStudy (pgreene008):

is that correct?

OpenStudy (pgreene008):

@wio

OpenStudy (anonymous):

Let me see

OpenStudy (pgreene008):

ok @wio

OpenStudy (anonymous):

Yes, it is correct

OpenStudy (pgreene008):

thanks

OpenStudy (pgreene008):

can you help me with this one? : Evaluate the expression. r = <9, -7, -1>, v = <2, 2, -2>, w = <-5, -2, 6> v ⋅ w @wio

OpenStudy (pgreene008):

i think its -26

OpenStudy (anonymous):

That is right

OpenStudy (pgreene008):

last one: Express the complex number in trigonometric form. -2 + 2 sqrt(3) i @wio

OpenStudy (pgreene008):

@Astrophysics

OpenStudy (anonymous):

You already know complex numbers and you were having trouble with dot products?

OpenStudy (dan815):

UdotV=|U||V|cosa

OpenStudy (anonymous):

Given \(z=x+iy\). We say \(z=re^{i\theta}\) where\[ r=\sqrt{x^2+y^2},\quad \theta = \arctan\left(\frac{y}{x}\right) \]

OpenStudy (dan815):

a=arccos(u . v/(sqrt(u . u +v . v)))

OpenStudy (pgreene008):

this is what i got: 4(cos5pi/6 + i sin 5pi/6)

OpenStudy (pgreene008):

@dan815 @wio

OpenStudy (astrophysics):